# AN ACCELERATED PRIMAL DUAL ALGORITHM WITH BACKTRACKING FOR DECENTRALIZED CONSTRAINED OPTIMIZATION

### v1

@@ 5. Parameter Choices
In the previous section, we considered a meta algorithm, i.e., given some arbitrary initial points $\mathbf{x}^{0}\in\mathop{\bf dom}\phi$, $\theta^{0}\in\mathcal{K}^{*}$, $\lambda^{0}\in\mathbb{R}^{n|\mathcal{E}|}$, step sizes $\gamma^{k}>0$, $\tau_{i}^{k},\sigma_{i}^{k}>0$ for $i\in\mathcal{N}$, and the momentum parameter $\eta^{k}>0$ for all $k\geq 0$, we set $\mathbf{x}^{-1}=\mathbf{x}^{0}$, $\theta^{-1}=\theta^{0}$, $\lambda^{-1}=\lambda^{0}$, and consider the iterations as in (3.10). For this generic framework, we established a duality gap result given in (4.15) assuming that there exists $\{t_{k}\}_{k\geq 0}\subset\mathbb{R}_{++}$ such that the algorithm parameters satisfy eq. 4.14; furthermore, we were able to bound the error term $P^{k}$ assuming that $\gamma^{k}\leq\Big(\sum_{i\in\mathcal{N}}d_{i}(\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}}{\alpha_{i}^{k+1}}+\frac{\eta^{k}}{\varsigma_{i}^{k}})\Big)^{-1}$. In this section, we will show that D-APDB displayed in algorithm 1 satisfies all these conditions stated above. Towards this goal, we first argue that D-APDB is well defined by showing that the agent-specific local backtracking conditions can be satisfied in finite backtracking iterations for all $i\in\mathcal{N}$.
@@
Next, we provide a sufficient condition on the stepsizes for (5.3) to hold. First, we provide some useful bounds on partial gradient differences. Given arbitrary $\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime}\in\mathop{\bf dom}\phi$ and $\theta^{\prime},\theta^{\prime\prime}\in\mathcal{K}^{*}\cap\mathcal{B}$. Let $\bar{L}_{x_{i}x_{i}}\triangleq L_{g_{i}}B_{i}$ for $i\in\mathcal{N}$. Thus, for any $i\in\mathcal{N}$ and $\lambda\in\mathbb{R}^{n|\mathcal{E}|}$, Assumption 3 implies that
@@
Then, for all $k\geq 0$, the test condition in Line 18 of D-APDB holds whenever at the time of checking it, $\tilde{\tau}_{i}^{k}$ satisfies the following two conditions:
@@
moreover, (5.6) is guaranteed to hold whenever $\tilde{\tau}_{i}^{k}\in(0,\hat{\tau}_{i}]$ where
$\hat{\tau}_{i}\triangleq\min\left\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{-L_{f_{i}}+\sqrt{L_{f_{i}}^{2}+4(1-(\delta+c))L_{g_{i}}^{2}B_{i}^{2}/c_{\beta}}}{2L_{g_{i}}^{2}B_{i}^{2}/c_{\beta}}},\ \frac{1}{C_{g_{i}}}\sqrt{\frac{{c_{\alpha}(1-\delta)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\zeta_{i}}}\right\}.$} (5.7)
Finally, one also has $\hat{\tau}_{i}\geq\min\Big\{\frac{1-(\delta+c)}{2L_{f_{i}}},~\frac{1}{L_{g_{i}}B_{i}}\sqrt{\frac{c_{\beta}(1-(\delta+c))}{2}},~\frac{1}{C_{g_{i}}}\sqrt{\frac{{c_{\alpha}(1-\delta)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\zeta_{i}}}\Big\}$.
@@
First, note that $\tilde{\Lambda}_{i}^{k}\leq\frac{L_{f_{i}}}{2}\|\tilde{x}_{i}^{k+1}-x_{i}^{k}\|^{2}$ as $\nabla f_{i}$ is Lipschitz with constant $L_{i}$ over $\mathop{\bf dom}\phi_{i}$. Morover, since $g_{i}$ and $\mathbf{J}{g_{i}}$ are Lipschitz on $\mathop{\bf dom}\phi_{i}$ with constants $C_{g_{i}}$ and $L_{g_{i}}$, respectively, and $\|\theta_{i}^{k}\|\leq B_{i}$, using (5.4) and (5.5), we get
@@
$\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\tilde{\tau}_{i}^{k}}{c_{\alpha}}}2C_{g_{i}}^{2}\|\tilde{\theta}_{i}^{k+1}-\theta_{i}^{k}\|^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\tilde{\tau}_{i}^{k}}{c_{\beta}}}L_{g_{i}}^{2}B_{i}^{2}\|\tilde{x}_{i}^{k+1}-x_{i}^{k}\|^{2};$
therefore, since $\tilde{\sigma}_{i}^{k}=\zeta_{i}\tilde{\tau}_{i}^{k}$, we can conclude that (5.3) holds if (5.6) holds.
@@
Given a candidate primal size $\tilde{\tau}_{i}^{k}\in(0,\tau_{i}^{k-1}]$, let candidate dual step size $\tilde{\sigma}_{i}^{k}$ and momentum parameter $\eta_{i}^{k}$ together with test function parameters $\tilde{\alpha}_{i}^{k+1},\tilde{\beta}_{i}^{k+1},\tilde{\varsigma}_{i}^{k+1}$ and $\alpha_{i}^{k},\beta_{i}^{k},\varsigma_{i}^{k}$ are set as in (5.1) and (5.2). The D-APDB parameters $(\tilde{\tau}_{i}^{k},\tilde{\sigma}_{i}^{k},\eta_{i}^{k})$ is called admissible if the test condition in Line 18 of D-APDB holds with these parameters.
In Theorem 5.1 we established that for each $k\geq 0$, the test condition in Line 18 of D-APDB after finite number of backtracking iterations. For iteration $k\geq 0$, let the candidate primal step size is $\tilde{\tau}_{i}^{k}$ when the condition in Line 18 holds for $i\in\mathcal{N}$; hence, $\eta_{i}^{k}=\frac{\tau_{i}^{k-1}}{\tilde{\tau}_{i}^{k}}$, which implies that the number of times node-$i$ employ backtracking (i.e., shrinks its primal step size) within the $k$-th iteration is equal to $n_{i}^{k}\triangleq\log_{1/\rho}\eta_{i}^{k}$. Thus, $\eta^{k}=\max_{i\in\mathcal{N}}\{\eta_{i}^{k}\}=\max_{i\in\mathcal{N}}\{\tau_{i}^{k-1}/\tilde{\tau}_{i}^{k}\}$ denotes the largest amount of contraction of primal step size among all the nodes.
@@
The accepted parameters $\tau_{i}^{k},\sigma_{i}^{k},\eta^{k},\alpha_{i}^{k+1},\beta_{i}^{k+1},\varsigma_{i}^{k+1}$ are set after the max consensus step $\eta^{k}=\max_{i\in\mathcal{N}}\{\eta_{i}^{k}\}$, where $\tau_{i}^{k}={\tau_{i}^{k-1}}/{\eta^{k}}$, $\sigma_{i}^{k}=\zeta_{i}\tau_{i}^{k}$, and $\alpha_{i}^{k+1}=c_{\alpha}/\tau_{i}^{k}$, $\beta_{i}^{k+1}=c_{\beta}/\tau_{i}^{k}$, $\varsigma_{i}^{k+1}=c_{\varsigma}/\tau_{i}^{k}$. Then, each node-$i$ computes $x_{i}^{k+1}$ and $\theta_{i}^{k+1}$ according to Line 32 and Line 33 of D-APDB, respectively, for $i\in\mathcal{N}$. It is essential to note that $\eta^{k}=1$ implies no backtracking is employed, i.e., $\eta_{i}^{k}=1$ for all $i\in\mathcal{N}$; hence, $\tau_{i}^{k}=\tilde{\tau}_{i}^{k}=\tau_{i}^{k-1}$ for all $i\in\mathcal{N}$, which also implies that $\tilde{x}_{i}^{k+1}={\rm prox}_{\tau_{i}^{k}\phi_{i}}\Big(x_{i}^{k}-\tau_{i}^{k}\big(\nabla f_{i}(x_{i}^{k})+p_{i}^{k}\big)\Big)$ and $\tilde{\theta}_{i}^{k+1}=\mathcal{P}_{\mathcal{K}_{i}^{*}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cap\mathcal{B}_{i}}}\!\left(\theta_{i}^{k}+\sigma_{i}^{k}\,g_{i}(x_{i}^{k+1})\right)$ – that is why we set $x_{i}^{k+1}=\tilde{x}_{i}^{k+1}$ and $\theta_{i}^{k+1}=\tilde{\theta}_{i}^{k+1}$ in Line 35 and Line 36 of D-APDB for the scenario $\eta^{k}=1$.
Below we collect all the parameter conditions stated so far within the hypothesis of Lemmas 4.2, 4.3 and 5.1 together with (5.1) and (5.2). All of these required conditions are formally stated below in Condition 1.
@@
Suppose Assumptions 2, 3 and 1 hold, and $\delta,c_{\alpha},c_{\beta},c_{\varsigma}>0$ are given that satisfy $\delta+c<1$, where $c\triangleq c_{\alpha}+c_{\beta}+c_{\varsigma}$. For each $i\in\mathcal{N}$, let $x_{i}^{0}\in\mathop{\bf dom}\phi_{i},\theta_{i}^{0}\in\mathcal{K}_{i}^{*}$ be the initial primal-dual variables, $B_{i}$ denotes a dual bound satisfying Assumption 4, let $\bar{\tau}_{i},\zeta_{i}>0$ and $\rho\in(0,1)$ denote the step size parameters of node $i$ and the contraction coefficient. Then, the primal-dual step-size sequence $\{(\tau_{i}^{k},\sigma_{i}^{k})\}_{k\geq 0}\subset\mathbb{R}_{++}\times\mathbb{R}_{++}$ for $i\in\mathcal{N}$ and the momentum parameter sequence $\{(\gamma^{k},\eta^{k})\}_{k\geq 0}\subset\mathbb{R}_{++}\times\mathbb{R}_{++}$ generated according to D-APDB are acceptable with respect to Condition 1.
@@ 6. Convergence guarantees for D-APDB and the Proof of Main Results
Now we are ready to establish convergence guarantees for D-APDB. We first analyze how we can bound the duality gap for D-APDB primal-dual iterate sequence through carefully controlling the difference between the admissible parameters $(\tilde{\tau}_{i}^{k},\tilde{\sigma}_{i}^{k},\eta_{i}^{k})$ according to Definition 5.2 and the updated parameters after max consensus among the nodes $(\tau_{i}^{k},\sigma_{i}^{k},\eta^{k})$.
@@
Since $t_{0}=1$, from the recursive definition of $\{t_{k}\}$, it follows that $t_{k}=\Big(\Pi_{\ell=1}^{k}\eta^{\ell}\Big)^{-1}$ for $k\geq 1$. Moreover, since $\tau_{i}^{k}=\tau_{i}^{k-1}/\eta^{k}$ for $i\in\mathcal{N}$ and $k\geq 1$, we have $\Pi_{\ell=1}^{k}\eta^{\ell}=\frac{\tau_{i}^{0}}{\tau_{i}^{k}}$ for $i\in\mathcal{N}$ and $k\geq 1$. Because $\tau_{i}^{k}=\tau_{i}^{k-1}/\eta^{k}$ for $i\in\mathcal{N}$, whenever at least one node shrinks its step size, all the other nodes must shrink theirs as well. Moreover, for all $k\geq 0$ such that $\eta^{k}>1$, we have $\log_{1/\rho}\eta^{k}\geq 1$, i.e., every node must shrink their step size from $\tau_{i}^{-1}$ to $\tau_{i}^{k}$ by using at least one contraction by $\rho\in(0,1)$. On the other hand, since $\{\tau_{i}^{k}\}_{k\geq 0}$ is nonincreasing sequences for all $i\in\mathcal{N}$, it follows from Theorem 5.1 that whenever $\tau_{i}^{k^{*}-1}\leq\hat{\tau}_{i}$ for all $i\in\mathcal{N}$ for some $k^{*}\geq 0$, the test condition in Line 18 of D-APDB is guaranteed to hold for all $k\geq k^{*}$, i.e., $\eta^{k}=1$ for all $k\geq k^{*}$. Therefore, we can conclude that $|\mathcal{I}|\leq\Big\lfloor\log_{1/\rho}\Big(\max_{i\in\mathcal{N}}\frac{\bar{\tau}_{i}}{\hat{\tau}_{i}}\Big)\Big\rfloor+1$; thus, $\tau_{i}^{k}\geq\bar{\tau}_{i}\rho^{|\mathcal{I}|}\geq\rho\bar{\tau}_{i}\min_{j\in\mathcal{N}}\frac{\hat{\tau}_{j}}{\bar{\tau}_{j}}$ for all $i\in\mathcal{N}$. Using our earlier observation $t_{k}=\frac{\tau_{i}^{k}}{\tau_{i}^{0}}$ for all $i\in\mathcal{N}$ and $k\geq 1$, we obtain the following lower bound: $t_{k}\geq\rho\min_{i\in\mathcal{N}}\frac{\hat{\tau}_{i}}{\bar{\tau}_{i}}$, which follows from $\tau_{i}^{0}\leq\bar{\tau}_{i}$ for $i\in\mathcal{N}$.
Theorem 6.3.
@@
$\sum_{i\in\mathcal{N}}\bar{P}_{i}^{k}\leq-\frac{\delta}{2}\sum_{i\in\mathcal{N}}\Big(\frac{1}{\tau_{i}^{k}}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}+\frac{1}{\sigma_{i}^{k}}\|\theta_{i}^{k+1}-\theta_{i}^{k}\|^{2}\Big)+\mathbf{1}_{\mathcal{I}}(k)\sum_{i\in\mathcal{N}}\Xi_{i},$ (6.1)
where $\mathcal{I}$ is given in Definition 6.1, $\bar{P}_{i}^{k}$ is defined in (4.17), $\Xi_{i}\triangleq 2\Big(L_{f_{i}}+\frac{\bar{\tau}_{i}}{c_{\beta}}B_{i}^{2}L_{g_{i}}^{2}\Big)D_{i}^{2}+4\frac{\bar{\tau}_{i}}{c_{\alpha}}C_{g_{i}}^{2}B_{i}^{2}$ for $i\in\mathcal{N}$, and $\mathbf{1}_{\mathcal{I}}(k)=1$ if $k\in\mathcal{I}$, equal to 0 otherwise. Moreover, it also holds that
where $\Xi\triangleq\Big(\Big\lfloor\log_{1/\rho}\Big(\max_{i\in\mathcal{N}}\frac{\bar{\tau}_{i}}{\hat{\tau}_{i}}\Big)\Big\rfloor+1\Big)\sum_{i\in\mathcal{N}}\Xi_{i}$.

### v2

@@ 5. Parameter Choices
In the previous section, we considered a meta algorithm, i.e., given some arbitrary initial points $\mathbf{x}^{0}\in\mathop{\bf dom}\phi$, $\theta^{0}\in\mathcal{K}^{*}$, $\lambda^{0}\in\mathbb{R}^{n|\mathcal{E}|}$, step sizes $\gamma^{k}>0$, $\tau_{i}^{k},\sigma_{i}^{k}>0$ for $i\in\mathcal{N}$, and the momentum parameter $\eta^{k}>0$ for all $k\geq 0$, we set $\mathbf{x}^{-1}=\mathbf{x}^{0}$, $\theta^{-1}=\theta^{0}$, $\lambda^{-1}=\lambda^{0}$, and consider the iterations as in (3.8). For this generic framework, we established a duality gap result given in (4.14) assuming that there exists $\{t_{k}\}_{k\geq 0}\subset\mathbb{R}_{++}$ such that the algorithm parameters satisfy (4.13); furthermore, we were able to bound the error term $P^{k}$ assuming that $\gamma^{k}\leq\Big(\sum_{i\in\mathcal{N}}d_{i}(\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}}{\alpha_{i}^{k+1}}+\frac{\eta^{k}}{\varsigma_{i}^{k}})\Big)^{-1}$. In this section, we will show that D-APDB displayed in Algorithm 1 satisfies all these conditions stated above. Towards this goal, we first argue that D-APDB is well defined by showing that the agent-specific local backtracking conditions can be satisfied in finite backtracking iterations for all $i\in\mathcal{N}$.
@@
Next, we provide a sufficient condition on the stepsizes for (5.3) to hold. First, we provide some useful bounds on partial gradient differences. Consider some arbitrary $\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime}\in\mathop{\bf dom}\phi$ and $\theta^{\prime},\theta^{\prime\prime}\in\mathcal{K}^{*}\cap\mathcal{B}$. Let $\bar{L}_{x_{i}x_{i}}\triangleq L_{g_{i}}B_{i}$ for $i\in\mathcal{N}$ such that $L_{g_{i}}>0$; thus, for any $\lambda\in\mathbb{R}^{n|\mathcal{E}|}$, Assumption 3 implies that
@@
Then, for all $k\geq 0$, the test condition in Line 18 of D-APDB holds whenever at the time of checking it, $\tilde{\tau}_{i}^{k}$ satisfies the following two conditions, with the convention $L_{g_{i}}B_{i}=0$ for $L_{g_{i}}=0,$
@@
moreover, (5.6) is guaranteed to hold whenever $\tilde{\tau}_{i}^{k}\in(0,\hat{\tau}_{i}]$, where
$\hat{\tau}_{i}\triangleq\min\left\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{-L_{f_{i}}+\sqrt{L_{f_{i}}^{2}+4(1-(\delta+c))L_{g_{i}}^{2}B_{i}^{2}/c_{\beta}}}{2L_{g_{i}}^{2}B_{i}^{2}/c_{\beta}}},\ \frac{1}{C_{g_{i}}}\sqrt{\frac{{c_{\alpha}(1-\delta)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\zeta_{i}}}\right\}$} (5.7)
for $i\in\mathcal{N}$ such that $L_{g_{i}}>0$, and $\hat{\tau}_{i}\triangleq\min\{\frac{1-(\delta+c)}{2L_{f_{i}}},\frac{1}{C_{g_{i}}}\sqrt{\frac{c_{\alpha}(1-\delta)}{2\zeta_{i}}}\}$ for $i\in\mathcal{N}$ such that $L_{g_{i}}=0$. Finally, one also has55With the convention that $1/0=\infty$ when $L_{g_{i}}B_{i}=0$ for $\in\mathcal{N}$ such that $L_{g_{i}}=0$.
$\hat{\tau}_{i}\geq\min\Big\{\frac{1-(\delta+c)}{2L_{f_{i}}},~\frac{1}{L_{g_{i}}B_{i}}\sqrt{\frac{c_{\beta}(1-(\delta+c))}{2}},~\frac{1}{C_{g_{i}}}\sqrt{\frac{{c_{\alpha}(1-\delta)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\zeta_{i}}}\Big\},\quad\forall i\in\mathcal{N}.$
@@
First, note that $\tilde{\Lambda}_{i}^{k}\leq\frac{L_{f_{i}}}{2}\|\tilde{x}_{i}^{k+1}-x_{i}^{k}\|^{2}$ as $\nabla f_{i}$ is Lipschitz with constant $L_{f_{i}}$ over $\mathop{\bf dom}\phi_{i}$. Morover, since $g_{i}$ and $\mathbf{J}{g_{i}}$ are Lipschitz on $\mathop{\bf dom}\phi_{i}$ with constants $C_{g_{i}}$ and $L_{g_{i}}$, respectively, and $\|\theta_{i}^{k}\|\leq B_{i}$, using (5.4) and (5.5), we get
@@
$\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\tilde{\tau}_{i}^{k}}{c_{\alpha}}}2C_{g_{i}}^{2}\|\tilde{\theta}_{i}^{k+1}-\theta_{i}^{k}\|^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\tilde{\tau}_{i}^{k}}{c_{\beta}}}L_{g_{i}}^{2}B_{i}^{2}\|\tilde{x}_{i}^{k+1}-x_{i}^{k}\|^{2}.$
Since $\tilde{\sigma}_{i}^{k}=\zeta_{i}\tilde{\tau}_{i}^{k}$, we can conclude that (5.3) holds if (5.6) holds.
@@
Given a candidate primal step size $\tilde{\tau}_{i}^{k}\in(0,\tau_{i}^{k-1}]$, let candidate dual step size $\tilde{\sigma}_{i}^{k}$ and momentum parameter $\eta_{i}^{k}$ together with test function parameters $\tilde{\alpha}_{i}^{k+1},\tilde{\beta}_{i}^{k+1},\tilde{\varsigma}_{i}^{k+1}$ and $\alpha_{i}^{k},\beta_{i}^{k},\varsigma_{i}^{k}$ are set as in (5.1) and (5.2). The D-APDB parameters $(\tilde{\tau}_{i}^{k},\tilde{\sigma}_{i}^{k},\eta_{i}^{k})$ are called admissible if the test condition in Line 18 of D-APDB holds with these parameters.
In Theorem 5.1 we established that for each $k\geq 0$, the test condition in Line 18 of D-APDB holds after a finite number of backtracking iterations. For $k\geq 0$, let the candidate primal step size is $\tilde{\tau}_{i}^{k}$ when the condition in Line 18 holds for $i\in\mathcal{N}$; hence, $\eta_{i}^{k}=\frac{\tau_{i}^{k-1}}{\tilde{\tau}_{i}^{k}}$, which implies that the number of times node-$i$ employs backtracking (i.e., shrinks its primal step size) within the $k$-th iteration is equal to $n_{i}^{k}\triangleq\log_{1/\rho}\eta_{i}^{k}$. Thus, $\eta^{k}=\max_{i\in\mathcal{N}}\{\eta_{i}^{k}\}=\max_{i\in\mathcal{N}}\{\tau_{i}^{k-1}/\tilde{\tau}_{i}^{k}\}$ denotes the largest amount of contraction of primal step size among all the nodes.
@@
The accepted parameters $\tau_{i}^{k},\sigma_{i}^{k},\eta^{k},\alpha_{i}^{k+1},\beta_{i}^{k+1},\varsigma_{i}^{k+1}$ are set after the max-consensus step $\eta^{k}=\max_{i\in\mathcal{N}}\{\eta_{i}^{k}\}$, where $\tau_{i}^{k}={\tau_{i}^{k-1}}/{\eta^{k}}$, $\sigma_{i}^{k}=\zeta_{i}\tau_{i}^{k}$, and $\alpha_{i}^{k+1}=c_{\alpha}/\tau_{i}^{k}$, $\beta_{i}^{k+1}=c_{\beta}/\tau_{i}^{k}$, $\varsigma_{i}^{k+1}=c_{\varsigma}/\tau_{i}^{k}$. Then, each node-$i$ computes $x_{i}^{k+1}$ and $\theta_{i}^{k+1}$ according to Line 32 and Line 33 of D-APDB, respectively, for $i\in\mathcal{N}$. It is essential to note that $\eta^{k}=1$ implies no backtracking is employed, i.e., $\eta_{i}^{k}=1$ for all $i\in\mathcal{N}$; hence, $\tau_{i}^{k}=\tilde{\tau}_{i}^{k}=\tau_{i}^{k-1}$ for all $i\in\mathcal{N}$, which also implies that $\tilde{x}_{i}^{k+1}={\rm prox}_{\tau_{i}^{k}\phi_{i}}\Big(x_{i}^{k}-\tau_{i}^{k}\big(\nabla f_{i}(x_{i}^{k})+p_{i}^{k}\big)\Big)$ and $\tilde{\theta}_{i}^{k+1}=\mathcal{P}_{\mathcal{K}_{i}^{*}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cap\mathcal{B}_{i}}}\!\left(\theta_{i}^{k}+\sigma_{i}^{k}\,g_{i}(x_{i}^{k+1})\right)$ – that is why we set $x_{i}^{k+1}=\tilde{x}_{i}^{k+1}$ and $\theta_{i}^{k+1}=\tilde{\theta}_{i}^{k+1}$ in Line 35 and Line 36 of D-APDB for the scenario $\eta^{k}=1$.
Below we collect all the parameter conditions stated so far within the hypothesis of Lemmas 4.3, 4.4 and 5.1 together with (5.1) and (5.2). All of these required conditions are formally stated below in Condition 1.
@@
Suppose Assumptions 2, 3 and 1 hold, and $\delta,c_{\alpha},c_{\beta},c_{\varsigma}>0$ are given that satisfy $\delta+c<1$, where $c\triangleq c_{\alpha}+c_{\beta}+c_{\varsigma}$. For each $i\in\mathcal{N}$, let $x_{i}^{0}\in\mathop{\bf dom}\phi_{i},\theta_{i}^{0}\in\mathcal{K}_{i}^{*}$ be the initial primal-dual variables, and let $\bar{\tau}_{i},\zeta_{i}>0$ and $\rho\in(0,1)$ denote the step size parameters of node $i$ and the contraction coefficient. Then, the primal-dual step-size sequence $\{(\tau_{i}^{k},\sigma_{i}^{k})\}_{k\geq 0}\subset\mathbb{R}_{++}\times\mathbb{R}_{++}$ for $i\in\mathcal{N}$ and the momentum parameter sequence $\{(\gamma^{k},\eta^{k})\}_{k\geq 0}\subset\mathbb{R}_{++}\times\mathbb{R}_{++}$ generated according to D-APDB are acceptable with respect to Condition 1.
@@ 6. Convergence guarantees for D-APDB and the Proof of Main Results
Now we are ready to establish convergence guarantees for D-APDB. We first analyze how we can bound the duality gap for D-APDB primal-dual iterate sequence through carefully controlling the difference between the admissible parameters $(\tilde{\tau}_{i}^{k},\tilde{\sigma}_{i}^{k},\eta_{i}^{k})$ according to Definition 5.2 and the updated parameters after computing max-consensus among the nodes $(\tau_{i}^{k},\sigma_{i}^{k},\eta^{k})$.
@@
Since $t_{0}=1$, from the recursive definition of $\{t_{k}\}$, it follows that $t_{k}=\Big(\Pi_{\ell=1}^{k}\eta^{\ell}\Big)^{-1}$ for $k\geq 1$. Moreover, for $i\in\mathcal{N}$ and $k\geq 1$, since $\tau_{i}^{k}=\tau_{i}^{k-1}/\eta^{k}$, we have $\Pi_{\ell=1}^{k}\eta^{\ell}=\frac{\tau_{i}^{0}}{\tau_{i}^{k}}$; hence, $t_{k}=\frac{\tau_{i}^{k}}{\tau_{i}^{0}}$. Due to $\tau_{i}^{k}=\tau_{i}^{k-1}/\eta^{k}$ for $i\in\mathcal{N}$, whenever at least one node shrinks its step size, all the other nodes must shrink theirs as well. Therefore, for all $k\geq 0$ such that $\eta^{k}>1$, we have $\log_{1/\rho}\eta^{k}\geq 1$, i.e., every node must shrink their step size from $\tau_{i}^{k-1}$ to $\tau_{i}^{k}$ by using at least one contraction by $\rho\in(0,1)$. On the other hand, since $\{\tau_{i}^{k}\}_{k\geq 0}$ is nonincreasing sequences for all $i\in\mathcal{N}$, it follows from Theorem 5.1 that whenever $\tau_{i}^{k^{*}-1}\leq\hat{\tau}_{i}$ for all $i\in\mathcal{N}$ for some $k^{*}\geq 0$, the test condition in Line 18 of D-APDB is guaranteed to hold for all $k\geq k^{*}$, i.e., $\eta^{k}=1$ for all $k\geq k^{*}$. Therefore, we can conclude that $|\mathcal{I}|\leq\Big\lfloor\log_{1/\rho}\Big(\max_{i\in\mathcal{N}}\frac{\bar{\tau}_{i}}{\hat{\tau}_{i}}\Big)\Big\rfloor+1$; thus, $\tau_{i}^{k}\geq\bar{\tau}_{i}\rho^{|\mathcal{I}|}\geq\rho\bar{\tau}_{i}\min_{j\in\mathcal{N}}\frac{\hat{\tau}_{j}}{\bar{\tau}_{j}}$ for all $i\in\mathcal{N}$. Using our earlier observation $t_{k}=\frac{\tau_{i}^{k}}{\tau_{i}^{0}}$ for all $i\in\mathcal{N}$ and $k\geq 1$, we obtain the following lower bound: $t_{k}\geq\rho\min_{i\in\mathcal{N}}\frac{\hat{\tau}_{i}}{\bar{\tau}_{i}}$, which follows from $\tau_{i}^{0}\leq\bar{\tau}_{i}$ for $i\in\mathcal{N}$.
Lemma 6.3.
@@
$\sum_{i\in\mathcal{N}}\bar{P}_{i}^{k}\leq-\frac{\delta}{2}\sum_{i\in\mathcal{N}}\Big(\frac{1}{\tau_{i}^{k}}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}+\frac{1}{\sigma_{i}^{k}}\|\theta_{i}^{k+1}-\theta_{i}^{k}\|^{2}\Big)+\mathbf{1}_{\mathcal{I}}(k)\sum_{i\in\mathcal{N}}\Xi_{i}^{k},$ (6.1)
where $\mathcal{I}$ is given in Definition 6.1, $\bar{P}_{i}^{k}$ is defined in (4.16), and $\mathbf{1}_{\mathcal{I}}(k)=1$ if $k\in\mathcal{I}$, equal to 0 otherwise; moreover, for $i\in\mathcal{N}$, for the case $L_{g_{i}}>0$, $\Xi_{i}^{k}\triangleq 2\Big(L_{f_{i}}+\frac{\bar{\tau}_{i}}{c_{\beta}}B_{i}^{2}L_{g_{i}}^{2}\Big)D_{i}^{2}+4\frac{\bar{\tau}_{i}}{c_{\alpha}}C_{g_{i}}^{2}B_{i}^{2}$, and for the case $L_{g_{i}}=0$, $\Xi_{i}^{k}=2L_{f_{i}}D_{i}^{2}+\frac{\bar{\tau}_{i}}{c_{\alpha}}C_{g_{i}}^{2}\|{\theta}_{i}^{k+1}-\theta_{i}^{k}\|^{2}$ for $k\geq 0$.